OpenStudy (anonymous):

how many gallons of water must be evaporated from 36 gallons of a 10% salt solution so that the resulting solution 12% salt solution

OpenStudy (anonymous):

how to do this Mr. Sourwing

OpenStudy (anonymous):

0.9(36) - x = 0.88(36-x) x = 6

OpenStudy (anonymous):

do u see why?

OpenStudy (anonymous):

yes thank you